How many positive integers less than 100 is not a factor of 2,3 and 5?


1
Expert's answer
2022-01-06T12:09:34-0500

Let "A" denote the set of positive integers less than 100 divisible 2.

Let "B" denote the set of positive integers less than 100 divisible 3.

Let "F" denote the set of positive integers less than 100 divisible 5.

Then


"N(A)=49, N(B)=33, N(C)=19,"

"N(A\cap B)=16, N(A\cap F)=9, N(B\cap F)=6,"

"N(A\cap B\cap F)=3"

"N(A\cup B\cup F)=N(A)+N(B)+N(F)"

"-N(A\cap B)-N(A\cap F)-N(B\cap F)"

"+N(A\cap B\cap F)"

"=49+33+19-16-9-6+3=73"

The number of positive integers less than 100, which are not divisible by 2, 3 or 5, is


"99-73=26"